OpenStudy (anonymous):

Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31 A. 5 B. 3 C. 2 D. 4

OpenStudy (anonymous):

By adding the two equations we get the x - value. Afterwards we simply substitute the computed x-value to any of the two given equations obtain the y-value. 3x - 5y = -16 (eq. 1) 2x + 5y = 31 (eq. 2) 5x = 15 x = 3 Let's try substituting x = 3 to the first eq. 1 to obtain y-value 3(3) - 5y = -16 9 - 5y = -16 -5y = -25 y = 5. (y-value is 5) You will get the same y-value if you insert x = 3 to eq. 2

OpenStudy (anonymous):

thank you